OpenStudy (lavenderandsage):

Find the odds against making a basket if you have made 17 out of 25 shots.

OpenStudy (kl0723):

25=100% 17= x% after cross multiplying you get that 17 is about 68% sucess rate, then there is a 32% chance that you won't make the next shot based on your previous performance

OpenStudy (kl0723):

25x = 100(17) 25x = 1700 x = 1700/25 x = 68% = sucess rate, meaning 17 shots are 68% of 25 shots that you attempted so the difference are the shots you've missed

OpenStudy (kl0723):

which is 32%

OpenStudy (lavenderandsage):

thank you so much!!!

OpenStudy (sshayer):

probability of odds against=\[\frac{ 25-17 }{ 25 }=\frac{ 8 }{ 25 }\]